A University found that 20% of its students withdraw without completing the first semester courses. Assume that 6 students have registered for the course this semester. To compute the following questions use binomial probability distribution.

a.      Compute the probability that 2 or fewer will withdraw?

b.      Compute the probability that exactly 4 will withdraw?

c.       Compute the expected number of withdrawals?

d.     What is the variance of the number of withdrawals?

e.     What is the standard deviation of the number of withdrawals?



1
Expert's answer
2021-03-24T13:50:39-0400

a. "p = 0.2 \Rightarrow q = 1 - p = 0.8"

Then the probabilities that 0,1 or 2 will withdraw are respectively equal

"{P_6}(0) = {q^6} = {0.8^6} = {\rm{0}}{\rm{.262144}}"

"{P_6}(1) = C_6^1p{q^5} = 6 \cdot 0.2 \cdot {0.8^5} = {\rm{0}}{\rm{.393216}}"

"{P_6}(2) = C_6^2{p^2}{q^4} = 15 \cdot {0.2^2} \cdot {0.8^4} = {\rm{0}}{\rm{.24576}}"

Then the required probability is

"P(X \le 2) = {P_6}(0) + {P_6}(1) + {P_6}(2) = 0.262144 + 0.393216 + 0.24576 = 0.90112"

Answer: "P(X \le 2) = 0.90112"

b. We have

"{P_6}(4) = C_6^4{p^4}{q^2} = 15 \cdot {0.2^4} \cdot {0.8^2} = {\rm{0}}{\rm{.01536}}"

Answer: "{P_6}(4) = {\rm{0}}{\rm{.01536}}"

c. The expected number is equal to the expected value. Then

"M(x) = np = 6 \cdot 0.2 = 1.2"

Answer: "M(x) = 1.2"

d. The variance of the number of withdrawals is "D(x) = npq = 6 \cdot 0.2 \cdot 0.8 = {\rm{0}}{\rm{.96}}".

Answer: 0.96

e. The standard deviation of the number of withdrawals is "\sigma (x) = \sqrt {D(x)} = \sqrt {0.96} \approx 0.98".

Answer: "\sigma (x) \approx 0.98"